The manager of a store selling tea plans to mix a more expensive tea that costs $ 9 per pound with a less expensive tea that costs $ 3 per pound to create a 70 pound blend that will sell for $ 4.80 per pound. How many pounds of each type of tea are required?
Answer by addingup(3677)

Let the amount of blend that costs 9 be x. Since we want a total quantity of 70, then the amount of $3 blend will be 70-x
:
9x+3(70-x) = 4.80(70)
9x+210-3x = 336
6x = 126
x = 21 He will need 21 pounds of the $9 tea and
70-21 = 49 pounds of the $3 tea to get 70 pounds at $4.80
:
John